OpenStudy (anonymous):

Solve the equation: x2 - 12x = -36

OpenStudy (saifoo.khan):

Quadratic formula. Safe and easy method. ;)

OpenStudy (anonymous):

x2-12x+36=0 x2-6x-6x+36=0 x(x-6)-6(x-6)=0 (x-6)^2=0 x=6...

OpenStudy (anonymous):

or factor it.... x2 - 12x + 36